Find the cost per pound of a coffee mixture made from 7 lb of coffee that costs $9.50 per pound and
3 lb of coffee that costs $8.10 per pound.

Answer by stanbon(75887)

Find the cost per pound of a coffee mixture made from 7 lb of coffee that costs $9.50 per pound and 3 lb of coffee that costs $8.10 per pound.
------
Total weight = 7+3 = 10 lb
Total value = 7*9.50 + 3*8.10 = 90.8 dollars
-----
Cost per pound = 90.8/10 = $9.08
=============
cheers,
Stan H.
---------------